Hello again! thank you very much for looking at this question


After watching a number of children playing games at a video arcade, a statistics practitioner estimated the following probability distribution of x, the number of games per visit:

x 1 2 3 4 5 6 7
P(x) 0.05 0.15 0.15 0.25 0.2 0.1 0.1

- Determine the probability P(1 THANK YOU AGAIN! im really stressed out right now i missed my stats class and your the only ones who can help me! i really appreciate this thank you!

Answer by jim_thompson5910(35256)

To find P(1 < x ≤ 4), simply add the probabilities from x=2 to x=4 (note: we are NOT including 1)

P(1 < x ≤ 4) = P(x=2) + P(x=3) + P(x=4)


P(1 < x ≤ 4) = 0.15+0.15+0.25


P(1 < x ≤ 4) = 0.55


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim